OpenStudy (anonymous):

solve log base 8 of 64=x

OpenStudy (anonymous):

do you know log rules? ... which rule can be applied here? :)

OpenStudy (anonymous):

\[\LARGE \log_a(b)=c\quad \Longrightarrow b=a^c\]

OpenStudy (anonymous):

log base 8 of 64 = log base 8 of 8^2 applying the rule log base a of b = c = a^c you get 2.

OpenStudy (anonymous):

\[\LARGE \log_{8}(64)=x\] \[\LARGE 64=8^x\] \[\LARGE 8^2=8^x\quad \iff 2=x\]

OpenStudy (anonymous):

The logarithm of a number is the power to which the base must be raised to equal the number. Because 8^2=64, log(base 8)64 = 2

jhonyy9 (jhonyy9):

x = xlog8(8) xlog8(8)=log8(64)=log8(8^2)=2log8(8) than x=2